OpenStudy (anonymous):

need help with this questionm, its improper integrals, and the question is integral from -infinity to infinity, xdx/(x^2+4)^3/2

OpenStudy (irishboy123):

try subbing out the x^2 + 4 and see what happens.

OpenStudy (anonymous):

this is Calc 2 you know

OpenStudy (irishboy123):

not sure what calc 2 means. this is a straight integral, but i thought a sub might make it easier if you do not see the direct approach. the limits of integration cannot be the problem, AFAICS. i will leave you to i unless there is something that you think i could add.

OpenStudy (dumbcow):

after making substitution ----> u = x^2 + 4 ----> du = 2x dx you have \[\frac{1}{2} \int\limits_{-\infty}^{\infty} \frac{du}{u^{3/2}} = -\frac{1}{\sqrt{u}} = -\frac{1}{\sqrt{x^2 +4}}\] Now apply limits, note that lim 1/infty = 0